OpenStudy (anonymous):

What is the Derivative of 4/x^3 ??

OpenStudy (luigi0210):

\[\frac{ 4 }{ x^3 }=4(x^(-3))\]

OpenStudy (luigi0210):

Just drop the power..

OpenStudy (anonymous):

4/x??

OpenStudy (primeralph):